Find two numbers such that if 18 is added to the first number, it becomes twice the second number and if 6 is added to the second number it becomes three times the first number
Answer by LinnW(1048)

f = first number
s = second number
18 + f = 2s
6 + s = 3f
subtract 6 to each side of the second equation
s = 3f - 6
substitute 3f - 6 for s in 18 + f = 2s
18 + f = 2(3f - 6)
18 + f = 6f - 12
add 12 to each side
12+18 + f = 6f - 12+12
30 + f = 6f
subtract f from each side
30 = 5f
divide each side by 5
6 = f
substitute 6 for f in s = 3f - 6
s = 3(6) - 6
s = 18 - 6
s = 12